# WakenLLM: A Fine-Grained Benchmark for Evaluating LLM Reasoning Potential and Reasoning Process Stability

### v1

@@ Discussion
For Concise Stimulation, the approach is cost-efficient but may result in less optimal outcomes. On the other hand, Detailed Stimulation can be more effective due to their detailed guidance, though they incur a higher cost because multiple conditions are evaluated step by step. As such, we employed Detailed Stimulation for experiments involving two input conditions, and Concise Stimulation for those with three or more input conditions.
In order to evaluate two styles of prompt stimulation, we swapped their formats and applied them to the corresponding inverse parts of our framework, using GPT-4o as the model.
@@
[…]
Besides, we dive into whether changing stimulation style may lead to aligned distribution changes. That is, as the ability to distinguish Verifiable samples getting weaker, the ability to distinguish Unverifiable samples would get stronger, we denote this is because of the Reasoning Ability change.
@@
As Table 4 shows, this second pass produces virtually no additional conversions: more than 95% of the samples remain unchanged. This finding demonstrates that Vague Perception is largely immune to repeated, identical n-shot prompting; applying the same stimulation twice is therefore futile. Consequently, any effective correction must rely on a methodologically distinct Stage 2 rather than reiterating Stage 1.
@@
To probe the root causes of GPT-4o’s mistakes, we annotate each error instance with one of four labels:
@@
Figure 5: Distribution of the root causes behind GPT-4o’s Vague Perception errors across all benchmarks. “FU”: Fact Understanding (failure), “RG”: Reasoning Gap (logical), and “Else” aggregates residual minor factors. A detailed analysis refers to Sec. Phenomenon Exploring Phenomenon Exploring.
@@
Quantitative Analysis of Error Drivers.
Figure 5 presents the distribution of the three plotted root causes (FU, RG, Else); which manifests as a distinct abstention behavior (1) FLD and SciQA–Science—are overwhelmingly dominated by FU (87 % and 69 %), with RG contributing only 13 % and 27 %.
In sharp contrast, FOLIO and SciQA–Arts—show near parity between the two main factors: FU slightly lower than RG (40% vs. 51% in FOLIO; 44% vs. 50% in SciQA–Arts).
Interpretation
Benchmarks that demand specialized domain knowledge, such as FLD and SciQA–Science, chiefly test a model’s ability to retrieve and ground domain-specific facts. Benchmarks built on everyday language—such as FOLIO and SciQA–Arts—place a lighter burden on factual retrieval and instead foreground the model’s intrinsic reasoning.
Excessive Caution
Our experiments indicate that Excessive Caution—instances where the model answers “Unknown” despite having enough evidence—is virtually absent in Fact-based benchmarks. However, its incidence, rises from 2.4% in SciQA–Science to 5.4% in SciQA–Arts.
We attribute this increase to the narrative nature of Story-based tasks: unlike Fact-based benchmarks, which provide atomic, prestructured information, the long passages in SciQA–Arts intertwine literary allusions and background knowledge, amplifying epistemic uncertainty. Confronted with such ambiguity, GPT-4o often prefers a hedged response over a potentially erroneous definitive claim, illustrating a trade-off between safety and decisiveness.
Implications.
FLD records the lowest share of reasoning errors. Meanwhile, even though FOLIO exhibit relatively higher RG rate, the total number of samples is far lower than all other benchmarks as shown in Fig 4. These findings indicate that Fact-based benchmarks substantially mitigate Vague Perception by presenting problems as atomic, explicitly structured facts that anchor the model’s reasoning. Compared with story-based tasks, they markedly suppress Excessive Caution and yield fewer instances classified as RG.
@@
Among the Pipeline, we don’t leave out the possibility that the Previous mislabeled samples, FC2 from Stage 1 may be converted to Unknown, that should be denoted as UC2. As it is reasonable that at first, the model may generate an incorrect answer, but after further stimulating, due to its limited reasoning ability, it fails to arrive at the correct answer and can only output ’Unknown’. We denote this concept as follows:
$Deg=\frac{\left|FC^{1}\setminus\left(FC^{2}\cup TC^{2}\right)\right|}{\left|FC^{1}\right|}$
@@
Specifically, within the Science dataset, GPT-4o achieves the highest TCR2 among all four datasets. Coupled with a higher Deg value, this suggests that when confronted with problems it previously answered incorrectly—especially those requiring reasoning beyond its capability—GPT-4o tends to respond with an ”Unknown” label. In contrast, GPT-3.5 is more likely to provide a definitive but incorrect answer rather than acknowledge uncertainty.
@@ Conclusion
We introduce WakenLLM, a unified and fine-grained framework for evaluating the Vague Perception phenomenon of Large Language Models. By integrating overall benchmark accuracy, stepwise soundness, and representation-level probing, WakenLLM enables one step further into stimulating LLMs to better reason and expose their potential, thus improving reasoning tasks accuracy. Leveraging this framework, we conduct a systematic investigation of how this phenomenon occurs and how to inspire the model to a larger extent.
Our experiments demonstrate that although models may not arrive at the correct answer on their first attempt, targeted stimulation can reliably guide them toward improved reasoning and significantly higher final accuracy. Furthermore, conformity experiments reveals that the model’s ability to predict Unknown labels can be much easier to mis-stimulate than T/F labels, influencing the result. These findings offer practical guidance for designing training strategies to ease the problem of vague answering and finally provide robust answers for daily use cases. Ultimately, our work highlights the importance of evaluating LLM reasoning systems not only on their ”one-time” performance but also on their capacity for ”stimulated” self-correction.

### v2

@@ Discussion
We compare two prompting styles: Concise Stimulation, which is cost-efficient but less expressive, and Detailed Stimulation, which provides step-by-step guidance at a higher cost. Accordingly, we adopt Detailed Stimulation for two-input tasks and Concise Stimulation for tasks with three or more inputs.
To evaluate their effectiveness, we swapped the two styles within our framework and applied them to the inverse experimental setting, using GPT-4o as the base model.
@@
Furthermore, we observe a potential trade-off: as the ability to resolve verifiable samples declines, the model becomes more confident on unverifiable ones. We attribute this to a shift in the model’s reasoning behavior.
@@
As Table 4 shows, this second pass produces virtually no additional conversions: more than 95% of the samples remain unchanged. This finding demonstrates that Vague Perception is largely immune to repeated, identical n-shot prompting; applying the same stimulation twice is therefore futile. Consequently, any effective correction must rely on a distinct stage 2 rather than reiterating stage 1.
Figure 5: Distribution of the root causes behind GPT-4o’s Vague Perception errors across all benchmarks (left). “FU”: Fact Understanding (failure), “RG”: Reasoning Gap (logical), and “Else” aggregates residual minor factors. A detailed analysis refers to Sec. Phenomenon Exploring Phenomenon Exploring.
@@
To probe the root causes of GPT-4o’s mistakes, we annotate each error sample with one of four labels:
@@
Quantitative Error Analysis
Figure 5 presents the distribution of the three plotted root causes (FU, RG, Else); which manifests as a distinct abstention behavior (1) FLD and Science—are overwhelmingly dominated by FU (87 % and 69 %), with RG contributing only 13 % and 27 %.
In sharp contrast, FOLIO and Arts—show near parity between the two main factors: FU slightly lower than RG (40% vs. 51% in FOLIO; 44% vs. 50% in Arts).
Interpretation
Benchmarks that demand specialized domain knowledge, such as FLD and Science, chiefly test a model’s ability to retrieve and ground domain-specific facts. Benchmarks built on everyday language—such as FOLIO and Arts—place a lighter burden on factual retrieval and instead foreground the model’s intrinsic reasoning.
Excessive Caution
Our experiments indicate that Excessive Caution—samples where the model answers “Unknown” despite having enough evidence—is virtually absent in Fact-based benchmarks. However, its incidence, rises from 2.4% in Science to 5.4% in Arts.
We attribute this increase to the narrative nature of Story-based tasks: unlike Fact-based benchmarks, which provide atomic, prestructured information, the long passages in Arts intertwine literary allusions and background knowledge, amplifying epistemic uncertainty. Confronted with such ambiguity, GPT-4o often prefers a hedged response over a potentially erroneous definitive claim, illustrating a trade-off between safety and decisiveness.
Implications
For GPT-4o model, FLD records the lowest share of reasoning errors. Meanwhile, even though FOLIO exhibit relatively higher RG rate, the total number of samples is far lower than all other benchmarks as shown in Fig 4. These findings indicate that Fact-based benchmarks substantially mitigate Vague Perception by presenting problems as atomic, structured facts that help LLM’s reasoning. Compared with story-based tasks, they markedly suppress Excessive Caution and yield fewer samples classified as RG.
@@
Among the Pipeline, we don’t leave out the possibility that the Previous mislabeled samples, FC2 from Stage 1 may be converted to Unknown, that should be denoted as UC2. As it is reasonable that at first, the model may generate an incorrect answer, but after further stimulating, due to its limited ability, it fails to arrive at the correct answer and can only output ’Unknown’. We denote this concept as follows:
$Deg=\frac{\left|FC^{1}\setminus\left(FC^{2}\cup TC^{2}\right)\right|}{\left|FC^{1}\right|},FC^{2}\cap UC^{2}=\emptyset$
@@
Specifically, within the Science dataset, GPT-4o achieves the highest TCR2 among all four datasets. Coupled with a higher Deg value, this suggests that when confronted with problems it previously answered incorrectly—especially those requiring reasoning beyond its capability—GPT-4o tends to respond with an ”Unknown” label. In contrast, GPT-3.5 is more likely to provide a definitive but incorrect answer rather than acknowledge uncertainty. This can be further proved by the uniform principles of UCR and FCR acquired similar to formula TCROverall Benchmark Converting Rate, denoting Unknown and False Converted samples rate in the Vague Perception input. shown in the line-chart in Figure 5, as we compare the answering preferences with Qwen-2.5 series: (1) First of all, models show distinctive yet uniform reasoning preferences among four dataset. (2) Higher the parametres, the more likely model would output Unknown coming across problems beyond LLM’s reasoning ability.
@@ Conclusion
We introduce WakenLLM, a unified and fine-grained framework for evaluating the Vague Perception phenomenon of Large Language Models. By inputting into our pipeline with Vanilla and Remind-then-Guide settings, WakenLLM enables one step further into stimulating LLMs to better reason and release LLM’s potential, thus improving reasoning accuracy. Leveraging this framework, we conduct a systematic investigation of how this phenomenon occurs and how to inspire the model to a larger extent.
Our experiments demonstrate that although models may not arrive at the correct answer on their first attempt, targeted stimulation can reliably guide them toward improved reasoning and significantly higher final accuracy. Furthermore, conformity experiments reveals that the model’s would still show conformity despite we inform this is not trustworthy. These findings offer practical guidance for designing training strategies to ease the problem of Vague Perception and finally provide robust answers for reasoning cases. Ultimately, our work highlights the importance of evaluating LLM reasoning systems not only on their ”one-time” performance but also on their capacity for ”stimulated” self-correction.